solve:
20 over x = 20 over x-4 + 4 over x
show work

Answer by stanbon(75887)

20/x = 20/(x-4) + 4/x
Multiply thru by x(x-4) to get:
20(x-4) = 20x +4(x-4)
20x-80 = 20x+4x-16
4x=-64
x=-16
Cheers,
Stan H.